OpenStudy (anonymous):

If f (x)=3x, g(x)=x-3, and h(x)= square root of x, find (fogoh)(25)

OpenStudy (shamil98):

\[f(x) = 3x , g(x) = x-3 , h(x) = \sqrt {x}\] Evaluate h(x) at 25, then input that into g(x) evaluate and then input that into f(x)

OpenStudy (anonymous):

It's f(g(h(x))) so you can either calculate h(25) and then for the number you will find replace on g(x) and so on. So h(25)= sqr(25)= 5 then g(5)=5-3=2 then f(2)=3*2 = 6 So (fogoh)(25)=6 ! You can replace the functions themselves if you want as well and then evaluate for 25

OpenStudy (shamil98):

\[\huge h(25) = 5\] \[\huge g(h(25)) = 5 - 3 = 2\] \[\huge f(g(h(25)) = 3(2) = 6\]